# Yong Ji1, Junye Li1 and Rui Yang∗2,3

### v3

@@ 4.1 Variational principle of level sets
Let $(X,d,f)$ be a t.d.s. with the specification property, and let $U\subset X$ be a nonempty open subset. Assume that $\varphi:M_{f}(X)\to\mathbb{R}$ is a continuous function satisfying Condition A. Then for any $\alpha\in L_{\varphi}$ and $S\in\{B,P\}$,
$\overline{\rm mdim}_{M}^{S}(f,K_{\alpha}^{\omega,U},d)=$ $\overline{\rm mdim}_{M}^{S}(f,K_{\alpha},d)$
$=$ $\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
[…]
@@
Define $G_{\mu}^{\omega,U}:=G_{\mu}\cap U\cap\left\{x\in X:\omega_{f}(x)=X\right\}$. By (3.17), there exists a constant $c>0$ such that
$\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{\mu}^{\omega,U},\epsilon)$
holds for all $\mu\in M_{f}(X)$ and sufficiently small $\epsilon>0$. Together with the fact that $G_{\mu}^{\omega,U}\subset K_{\alpha}^{\omega,U}$ holds for any $\mu\in M_{f}(\varphi,\alpha)$, we have
$\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,K_{\alpha}^{\omega,U},\epsilon)$
for all sufficiently small $\epsilon>0$. This implies that
$\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\leq\overline{\rm mdim}_{M}^{B}(f,K_{\alpha}^{\omega,U},d).$ (4.1)
Recall $G^{M_{f}(\varphi,\alpha)}=\{x\in X:V_{f}(x)\subset M_{f}(\varphi,\alpha)\}$. Fix $x\in X$. Since the statement $\varphi(\mu)=\alpha$ for all $\mu\in V_{f}(x)$ is equivalent to the statement $V_{f}(x)\subset M_{f}(\varphi,\alpha)$, it follows that $K_{\alpha}=G^{M_{f}(\varphi,\alpha)}$. By Lemma 3.2 $(1)$, we get
@@
By (4.1), (4.2), and the fact stated in Proposition 2.1 $\overline{\rm mdim}_{M}^{B}(f,Z,d)\leq\overline{\rm mdim}_{M}^{P}(f,Z,d)$ for any $Z\subset X$, we complete the proof. ∎
A direct consequence of Theorem 4.1 is the variational principles for the Bowen metric mean dimension of level sets, as derived by Backes and Rodrigues [BR23, Theorem A]. However, we are able to strengthen this result by demonstrating that the variational principle remains valid for the packing metric mean dimension of level sets [^3] Readers can turn to [O03, Tho09, ZC13] and references therein for more examples about the multifractal analysis of level sets using topological pressure..
Corollary 4.2.
Let $(X,d,f)$ be a t.d.s. with the specification property and $\psi$ be a continuous function on $X$. Let
$E_{\alpha}:=\{x\in X:\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\psi(f^{j}x)=\alpha\}$
denote the $\alpha$-level set of $\psi$. If $\alpha\in\mathbb{R}$ is a constant such that $E_{\alpha}\not=\emptyset$, then
$\overline{\rm mdim}_{M}^{B}(f,E_{\alpha},d)=\overline{\rm mdim}_{M}^{P}(f,E_{\alpha},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\psi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
[…]
Proof.
[…]
$K_{\alpha}=\{x\in X:\forall\mu\in V_{f}(x),\int\psi\mathop{}\!\mathrm{d}\mu=\alpha\}.$
Fix $x\in X$. Since the statement $\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\psi(f^{j}x)=\alpha$ is equivalent to the statement $V_{f}(x)\subset M_{f}(\psi,\alpha)$, this shows $K_{\alpha}=E_{\alpha}$. We obtain the desired result by Theorem 4.1. ∎
4.2 Linking metric mean dimension and mean Li-Yorke chaotic
In the present subsection, we establish a connection between the concepts of metric mean dimension and mean Li-Yorke chaos within the context of infinite entropy. We provide a quantitative description of the chaotic behavior exhibited by infinite entropy systems. Specifically, we demonstrate that the packing and upper capacity metric mean dimensions of the set of mean Li-Yorke pairs are equal to twice the metric mean dimension of the phase space, as stated in Theorem 4.3.
Given a t.d.s. $(X,d,f)$, this naturally induces a product system $(X\times X,d\times d,f\times f)$, where $X\times X$ is a compact space equipped with the product metric defined as
@@ 4.2 Linking metric mean dimension and mean Li-Yorke chaotic
and the mapping $f\times f:X\times X\rightarrow X\times X$, given by $(f\times f)(x,y):=(fx,fy)$, is continuous.
A pair $(x,y)\in X\times X$ is called a mean Li-Yorke pair if
@@
The set of all mean Li-Yorke pairs is denoted by ${\rm MLY}(f)$. A subset $S$ of $X$ is called a mean Li-Yorke set if any two distinct points in $S$ form a mean Li-Yorke pair. We say that the system $(X,f)$ is mean Li-Yorke chaotic if there is an uncountable mean Li-Yorke set. The well-known fact is that positive entropy implies mean Li-Yorke chaotic (cf.[Dow14]).
@@
Let $(X,d,f)$ be a t.d.s. with specification property. Then for $S\in\{UC,P\}$ we have
@@
It suffices to show the first equality since the second one is obtained similarly.
We claim that there exists $\mu_{0}\in M_{f\times f}(X\times X)$ such that $\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{0}(x,y)>0$.
Notice that the specification property implies that the system has positive topology entropy. Since positive entropy implies mean Li-Yorke chaotic (cf.[Dow14]), we can choose two distinct points of $x,y\in X$, a sequence $\{n_{i}\}$ of positive integers that converges to $\infty$ as $i\to\infty$, and $\theta>0$ (depending on $x,y$) such that
$\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}d(f^{j}x,f^{j}y)>\theta.$
Set $\mu_{n_{i}}=\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}\delta_{(f^{j}x,f^{j}y)}$. Then, in the weak-topology any accumulation point $\mu_{0}\in M_{f\times f}(X\times X)$ of $\{\mu_{n_{i}}\}$ satisfies the desired claim.
Now take $\nu\in M_{f}(X)$ arbitrarily and put $\mu_{1}:=\phi_{\ast}\nu=\nu\circ\phi^{-1}$, where $\phi:X\to X\times X$ is the diagonal map given by $\phi(x)=(x,x)$. It is easy to check that $\mu_{1}\in M_{f\times f}(X\times X)$, and $\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{1}=0$. For sufficiently small $\gamma>0$, let
$c_{\gamma}:=\min\{\frac{1}{\gamma},\overline{\rm mdim}_{M}^{UC}(f\times f,X\times X,d\times d)-\gamma\}.$
By Theorem 2.4, we pick a sequence of positive real numbers $\epsilon_{k}\to 0$ as $k$ goes to $\infty$ and a sequence of invariant measures $m_{k}\in M_{f\times f}(X\times X)$ such that for each $k$,
$\frac{1}{|\log\epsilon_{k}|}\inf_{{\rm diam}(\xi,d\times d)<\epsilon_{k}}h_{m_{k}}(f\times f,\xi)>c_{\gamma}.$ (4.3)
Let $K=\overline{{\rm Conv}}\{\mu_{0},\mu_{1},m_{k},k\geq 1\}$, i.e., the convex closure of the set $\{\mu_{0},\mu_{1}\}\cup\{m_{k}:k\geq 1\}$. Then $K$ is a closed convex subset of $M_{f\times f}(X\times X)$. It is easy to see that if two systems admit the specification property, so does the product system. By [PS07], we have $G_{K}\not=\emptyset$. For any $(x_{0},y_{0})\in G_{K}$, there are two sequences of integers $\{t_{n}^{0}\}_{n\in\mathbb{N}}$ and $\{t_{n}^{1}\}_{n\in\mathbb{N}}$ with $\mathcal{E}_{t_{n}^{0}}(x_{0},y_{0})\to\mu_{0}$, $\mathcal{E}_{t_{n}^{1}}(x_{0},y_{0})\to\mu_{1}$ as $n\to\infty$, respectively. Hence,
@@
This implies that $G_{K}\subset{\rm MLY}(f)$. Therefore, by Theorem 1.1, one has
@@
$=$ $\limsup_{\epsilon\to 0}\frac{\sup_{\nu\in K}\inf_{{\rm diam}(\eta,d\times d)<\epsilon}h_{\nu}(f\times f,\eta)}{|\log\epsilon|}$
$\geq$ $\limsup_{k\to\infty}\frac{\inf_{{\rm diam}(\xi,d\times d)<\epsilon_{k}}h_{m_{k}}(f\times f,\xi)}{|\log\epsilon_{k}|}\geq c_{\gamma}.$
Since $\gamma$ is arbitrary, one has
$\overline{\rm mdim}_{M}^{UC}(f\times f,X\times X,d\times d)\leq$ $\overline{\rm mdim}^{P}_{M}(f\times f,{\rm MLY}(f),d\times d)$
@@
$\leq$ $\overline{\rm mdim}_{M}^{UC}(f\times f,X\times X,d\times d),$
where we used the Proposition 2.1 for the second inequality. Together with $\overline{\rm mdim}_{M}^{UC}(f\times f,X\times X,d\times d)=2~{}\overline{\rm mdim}_{M}(f,X,d)$, we obtain the desired equality. ∎
Finally, we give an example to verify these results presented in Corollary 4.2 and Theorem 4.3.
Example 4.4.
Endow the product space $[0,1]^{\mathbb{Z}}$ with the product metric:
It is shown that the system $([0,1]^{\mathbb{Z}},d,\sigma)$ has the specification property (cf. [DGS76, Proposition 21.2]), and ${\rm{mdim}}_{M}(\sigma,[0,1]^{\mathbb{Z}},d)=1$ [LT18].
  • (1)
    Thus, by Theorem 4.3, we have
@@
  • (2)
    Let $\varphi:[0,1]^{\mathbb{Z}}\to\mathbb{R}$ be a continuous map defined by $\varphi((x_{n})_{n\in\mathbb{Z}})=x_{0}$. Then for any $\alpha\in[0,1]$, by Corollary 4.2 we have
    $\overline{\rm mdim}_{M}^{B}(\sigma,E_{\alpha},d)=\overline{\rm mdim}_{M}^{P}(\sigma,E_{\alpha},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{\sigma}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(\sigma,\xi),$
[…]
@@ 4.3 Generic points of full shift over compact metric space
In this subsection, based on the main results, we present some specific formulae for the metric mean dimensions of the set of generic points of full shifts over compact metric spaces.
We begin by reviewing some definitions and results for our statements. Let $(X,d,f)$ be a t.d.s., $\mu\in M_{f}(X)$, $\epsilon>0$ and $x\in X$. We define the Brin-Katok local entropy at $x$ as
$\overline{h}_{\mu}^{BK}(x,\epsilon)=\limsup_{n\to\infty}\frac{-\log\mu(B_{n}(x,\epsilon))}{n},$
and the Brin-Katok local $\epsilon$-entropy of $\mu$ as
$\overline{h}_{\mu}^{BK}(f,\epsilon):=\int\overline{h}_{\mu}^{BK}(x,\epsilon)\mathop{}\!\mathrm{d}\mu,$
respectively. Specially, if $\mu\in M_{f}^{e}(X)$, then $h_{\mu}^{BK}(x,\epsilon)=h_{\mu}^{BK}(f,\epsilon)$ holds for $\mu$-a.e. $x\in X$.
Example 4.5.
Let $K$ be a compact metric space with a metric $d$. Consider the left shift map $\sigma$ on $K^{\mathbb{Z}}$ again, where $K^{\mathbb{Z}}$ is endowed with the following product metric
$d^{\mathbb{Z}}\left((x_{n})_{n\in\mathbb{Z}},(y_{n})_{n\in\mathbb{Z}}\right)=\sum_{n\in\mathbb{Z}}\frac{d(x_{n},y_{n})}{2^{|n|}}.$
It is well-known that the system $(K^{\mathbb{Z}},\sigma)$ has the specification property. Moreover, it was proved in [VV17, Theorem 5] and [S22, Theorem 15] that,
$\overline{\rm mdim}_{M}(\sigma,K^{\mathbb{Z}},d^{\mathbb{Z}})=\overline{\rm dim}_{B}(K,d),$ (4.4)
where $\overline{\rm dim}_{B}(K,d)$ is the upper box dimension of $K$ under the metric $d$.
Then, by Corollary 1.5, the metric mean dimensions of the generic points of the full shift over compact metric space are fully determined by the boxing dimension of the underlying space and mean Rényi information dimension of invariant measures.
@@
Under the setting of Example 4.5, for any $\mu\in M_{\sigma}(K^{\mathbb{Z}})$ one has
@@
Next, we present two examples of full shifts with the maximal metric mean dimension measure.
Example 4.7.
Consider the Example 4.4 again. Let $\mathcal{L}$ be the Lebesgue measure on $[0,1]$, and define $\mu=\mathcal{L}^{\otimes\mathbb{Z}}$ as the product measure on $[0,1]^{\mathbb{Z}}$. Shi proved that $\lim_{\epsilon\to 0}\frac{h_{\mu}^{BK}(\sigma,\epsilon)}{|\log\epsilon|}=1$ [S22, Example 12]. Together with [YCZ25, Theorem 1.1], we have
$\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(\sigma,\xi)=1={\rm\overline{mdim}}_{M}(\sigma,[0,1]^{\mathbb{Z}},d).$
Then, by Corollary 1.5, we have
${\rm\overline{mdim}}_{M}^{B}(\sigma,G_{\mu},d)={\rm\overline{mdim}}_{M}^{P}(\sigma,G_{\mu},d)={\rm\overline{mdim}}_{M}^{UC}(\sigma,G_{\mu},d)=1.$
Example 4.8.
Endow $\mathbb{R}^{m}$ with the supremum norm $\|\cdot\|_{\mathbb{R}^{m}}$, and let $T_{i}:\mathbb{R}^{m}\to\mathbb{R}^{m}$ $1\leq i\leq l$ be contracting similitudes, that is,
$\|T_{i}x-T_{i}y\|_{\mathbb{R}^{m}}=r_{i}\|x-y\|_{\mathbb{R}^{m}}$
for some $r_{i}\in(0,1)$. One says that the iterated function system $\{T_{i}\}_{1\leq i\leq l}$ satisfies the Open Set Condition if there is a bounded open subset $V\subset\mathbb{R}^{m}$ such that $T_{i}(V)\subset V,\ \forall 1\leq i\leq l$, and $T_{i}(V)\cap T_{j}(V)=\emptyset,\ \text{for }i\neq j.$
Under the OSC condition, it is well known that there are (unique) compact subset $K$, which we call the self-similar set, and a measure $\mu$, which we call the self-similar measure such that
@@
    ${\rm dim}_{H}(K,d)={\rm dim}_{B}(K,d)=\alpha$, where $\alpha$ is the unique solution of equation
@@
    and ${\rm dim}_{H}$ denotes the Hausdorff dimension under the metric $d:=\|\cdot\|_{\mathbb{R}^{m}}$. $\alpha$ is called the similarity dimension of the IFS;
@@
    $\mu=\sum_{i=1}^{l}r_{i}^{\alpha}\cdot\mu\circ T_{i}^{-1}$;
  • (4)
    ${\rm supp}(\mu)=K$.
Many famous fractal sets, for instance, the middle thirds Cantor set, the Sierpinński triangle or gasket, are the attractors of iterated function systems satisfying the Open Set Condition.
Now take the attractor $K$ obtained above as in the Example 4.5. Then the product measure $\nu:=\mu^{\otimes\mathbb{Z}}$ on $K^{\mathbb{Z}}$ is invariant with respect to the left shift $\sigma$.
@@
Under the setting of Example 4.8, $\nu$ is the maximal metric mean dimension such that the three types of metric mean dimensions of $G_{\nu}$ coincide with the similarity dimension. i.e.,
$\overline{\rm mdim}_{M}^{S}(\sigma,G_{\nu},d^{\mathbb{Z}})=\alpha,$
where $S\in\{B,P,UC\}$
@@
Hutchinson [Hut81] showed that for any $z\in K$ and $\epsilon\in(0,1)$, one has
$\mu(B(z,\epsilon))\leq C\epsilon^{\alpha},$ (4.5)
where $C>0$ is a constant independent of the choice of the point  $z\in K$ and  $\epsilon$. For any $x=(x_{k})_{k\in\mathbb{Z}}\in K^{\mathbb{Z}}$ and $\epsilon\in(0,1)$, we define
[…]
Then the Bowen ball $B_{n}(x,\epsilon)$ is contained in $L_{n}(x,\epsilon)$ and $\nu(L_{n}(x,\epsilon))\leq C^{n}\epsilon^{n\alpha}$ by (4.5). Thus, one has
@@
which yields that $\overline{h}_{\nu}^{BK}(\sigma,\epsilon)\geq\alpha|\log\epsilon|-\log C.$ By the Shannon-McMillan-Breiman theorem, the following inequality is valid for any t.d.s. $(X,d,f)$:

### v4

@@ 4.1 Metric mean dimensions of level sets
Let $(X,d,f)$ be a TDS with the specification property and $\psi$ be a continuous function on $X$. If $\alpha\in\mathbb{R}$ is a constant such that $K_{\alpha}\not$$\emptyset$, then
$\overline{\rm mdim}_{M}^{B}(f,K_{\alpha},d)$ $=\overline{\rm mdim}_{M}^{P}(f,K_{\alpha},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\psi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
$\overline{\rm mdim}_{M}^{UC}(f,K_{\alpha},d)$ $=\overline{\rm mdim}_{M}(f,X,d).$
The corresponding result are also valid for the three types of lower metric mean dimensions by changing $\limsup_{\epsilon\to 0}$ into $\liminf_{\epsilon\to 0}$.
@@
If $K_{\alpha}\not$$\emptyset$, then $M_{f}(\psi,\alpha)$ is a non-empty convex set. For $x\in X$, the statement $\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\psi(f^{j}x)=\alpha$ is equivalent to the statement $V_{f}(x)\subset M_{f}(\psi,\alpha)$. Recall $G^{M_{f}(\varphi,\alpha)}=\{x\in X:V_{f}(x)\subset M_{f}(\varphi,\alpha)\}$. It follows that $K_{\alpha}=G^{M_{f}(\varphi,\alpha)}$. By Lemma 3.2 $(1)$, we get
@@
By inequality (3.17), there exists a constant $c>0$ such that for every $\mu\in M_{f}(X)$ and sufficiently small $\epsilon>0$,
$\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{\mu},\epsilon).$
Since $G_{\mu}\subset K_{\alpha}$ for any $\mu\in M_{f}(\varphi,\alpha)$, we have
$\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,K_{\alpha},\epsilon),$
which implies that
$\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\leq\overline{\rm mdim}_{M}^{B}(f,K_{\alpha},d).$ (4.2)
By (4.1), (4.2) and $\overline{\rm mdim}_{M}^{B}(f,K_{\alpha},d)\leq\overline{\rm mdim}_{M}^{P}(f,K_{\alpha},d)$, we get the first equalities.
Fix $\mu\in M_{f}(\psi,\alpha)$. Since the specification property implies $G_{\mu}\not$$\emptyset$ [PS07], we have $G_{\mu}\subset K_{\alpha}$. By Corollary 1.5, we obtain $\overline{\rm mdim}_{M}^{UC}(f,K_{\alpha},d)=\overline{\rm mdim}_{M}(f,X,d)$. ∎
Example 4.2.
Endow the product space $0,1]^{\mathbb{Z}}$ with the product metric:
It is shown that the system $([0,1]^{\mathbb{Z}},d,\sigma)$ has the specification property (cf. [DGS76, Proposition 21.2]), and ${\rm{mdim}}_{M}(\sigma,[0,1]^{\mathbb{Z}},d)=1$ [LT18, E. Example]. Let $\varphi:[0,1]^{\mathbb{Z}}\to\mathbb{R}$ be a continuous map given by $\varphi((x_{n})_{n\in\mathbb{Z}})=x_{0}$. Then for any $\alpha\in[0,1]$, by Theorem 4.1 we have
$\overline{\rm mdim}_{M}^{B}(\sigma,K_{\alpha},d)=\overline{\rm mdim}_{M}^{P}(\sigma,K_{\alpha},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{\sigma}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(\sigma,\xi),$
where $K_{\alpha}:=\{x\in[0,1]^{\mathbb{Z}}:\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}x_{j}=\alpha\}$, and $M_{\sigma}(\varphi,\alpha):=\{\mu\in M_{\sigma}([0,1]^{\mathbb{Z}}):\int\varphi d\mu=\alpha\}$. Furthermore, we have
$\overline{\rm mdim}_{M}^{UC}(\sigma,K_{\alpha},d)=1.$
4.2 Mean Li-Yorke chaos in infinite entropy systems
In this subsection, we provide a quantitative result involving the chaotic behavior of infinite entropy systems. Specifically, we prove that the (packing/upper capacity) metric mean dimension of the set of mean Li-Yorke pairs is equal to twice the metric mean dimension of the phase space.
Given a TDS $(X,d,f)$, this naturally induces a product system $(X\times X,d\times d,f\times f)$, where $X\times X$ is a compact space equipped with the product metric:
@@ 4.2 Mean Li-Yorke chaos in infinite entropy systems
and $(f\times f)(x,y):=(fx,fy)$. A pair $(x,y)\in X\times X$ is called a mean Li-Yorke pair if
@@
Namely, the orbits of $x$ and $y$ is sufficiently close to each other for sufficiently $n$, but they may be apart from each other for infinitely many time.
Denote by ${\rm MLY}(f)$ the set of all mean Li-Yorke pairs. A subset $S\subset X$ is called a mean Li-Yorke set if any two distinct points in $S$ form a mean Li-Yorke pair. The system $(X,f)$ is mean Li-Yorke chaotic if there is an uncountable mean Li-Yorke set. A well-known fact is that positive entropy systems are mean Li-Yorke chaotic [Dow14].
@@
Let $(X,d,f)$ be a TDS with the specification property. Then for $S\in\{UC,P\}$, we have
@@
It suffices to prove the first equality since the second one can be obtained in a similar manner.
We claim that there exists $\mu_{0}\in M_{f\times f}(X\times X)$ such that
$\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{0}(x,y)>0.$
Notice that the specification property implies that the system has positive topology entropy. Since positive entropy systems are mean Li-Yorke chaotic [Dow14], we choose two distinct points $x,y\in X$, a sequence $\{n_{i}\}$ of positive integers that converges to $\infty$ as $i\to\infty$, and $\theta>0$ (depending on $x,y$) such that $\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}d(f^{j}x,f^{j}y)>\theta.$ Put $\mu_{n_{i}}:=\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}\delta_{(f^{j}x,f^{j}y)}$. Then, in the weak-topology, any accumulation point $\mu_{0}\in M_{f\times f}(X\times X)$ of $\{\mu_{n_{i}}\}$ satisfies the desired claim.
Fix $\nu\in M_{f}(X)$ and put $\mu_{1}:=\nu\circ\phi^{-1}$, where $\phi:X\to X\times X$ is the diagonal map given by $\phi(x)=(x,x)$. It is easy to see that $\mu_{1}\in M_{f\times f}(X\times X)$ and $\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{1}=0$. Fix $\gamma>0$ and let
$c_{\gamma}:=\min\{\frac{1}{\gamma},\overline{\rm mdim}_{M}(f\times f,X\times X,d\times d)-\gamma\}.$
By Theorem 2.4, we choose a sequence of positive real numbers $\epsilon_{k}\to 0$ as $k\to\infty$ and a sequence $\{m_{k}\}$ of $f\times f$-invariant measures such that for every $k$,
$\frac{1}{|\log\epsilon_{k}|}\inf_{{\rm diam}(\xi,d\times d)<\epsilon_{k}}h_{m_{k}}(f\times f,\xi)>c_{\gamma}.$
Let
$K=\overline{{\rm co}}\{\mu_{0},\mu_{1},m_{k},k\geq 1\}.$
Then $K$ is a closed convex subset of $M_{f\times f}(X\times X)$. Since $(X\times X,d\times dmf\times f)$ has the specification property, by [PS07] we have $G_{K}\not$$\emptyset$. For any $(x_{0},y_{0})\in G_{K}$, there exists two sequences of integers $\{t_{n}^{0}\}$ and $\{t_{n}^{1}\}$ such that $\mathcal{E}_{t_{n}^{0}}(x_{0},y_{0})\to\mu_{0}$, $\mathcal{E}_{t_{n}^{1}}(x_{0},y_{0})\to\mu_{1}$ as $n\to\infty$, respectively. Hence,
@@
This yields that $G_{K}\subset{\rm MLY}(f)$. Therefore, by Theorem 1.1 one has
@@
$=$ $\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\nu\in K}\inf_{{\rm diam}(\eta,d\times d)<\epsilon}h_{\nu}(f\times f,\eta)$
$\geq$ $\limsup_{k\to\infty}\frac{1}{|\log\epsilon_{k}|}\inf_{{\rm diam}(\xi,d\times d)<\epsilon_{k}}h_{m_{k}}(f\times f,\xi)\geq c_{\gamma}.$
Letting $\gamma\to 0$, one has
$\overline{\rm mdim}_{M}(f\times f,X\times X,d\times d)\leq$ $\overline{\rm mdim}^{P}_{M}(f\times f,{\rm MLY}(f),d\times d)$
@@
$\leq$ $\overline{\rm mdim}_{M}(f\times f,X\times X,d\times d)$
$=$ $2~\overline{\rm mdim}_{M}(f,X,d).$
This completes the proof. ∎
For instance, under the setting of Example 4.2 we have
@@ 4.3 Generic points of full shift over compact metric space
For the full shifts over compact metric spaces, we present the precise formulae for the metric mean dimension of the set of generic points of invariant measures.
Let $(K,d_{K})$ be a compact metric space. Consider the left shift map $\sigma$ on $K^{\mathbb{Z}}$ again, where $K^{\mathbb{Z}}$ is endowed with the following product metric
$d\left((x_{n})_{n\in\mathbb{Z}},(y_{n})_{n\in\mathbb{Z}}\right)=\sum_{n\in\mathbb{Z}}\frac{d_{K}(x_{n},y_{n})}{2^{|n|}}.$
It is well-known that the system $(K^{\mathbb{Z}},d,\sigma)$ has the specification property. It was proved in [VV17, Theorem 5] that,
$\overline{\rm mdim}_{M}(\sigma,K^{\mathbb{Z}},d)$ $=\overline{\rm dim}_{B}(K,d_{K}),$ (4.3)
$\underline{\rm mdim}_{M}(\sigma,K^{\mathbb{Z}},d)$ $=\underline{\rm dim}_{B}(K,d_{K}),$ (4.4)
where $\overline{\rm dim}_{B}(K,d)$ and $\underline{\rm dim}_{B}(K,d)$ denote the upper and lower box dimensions of $K$ under the metric $d$, respectively.
By Corollary 1.5, for the full shift over compact metric spaces the metric mean dimensions of the set of generic points of invariant measures are fully determined by the box dimension of the underlying space and mean Rényi information dimension of invariant measures.
@@
Let $(K^{\mathbb{Z}},d,\sigma)$ be the full shift over $(K,d_{K})$. Then for any $\mu\in M_{\sigma}(K^{\mathbb{Z}})$,
@@
Next, for certain full shift over self-similar set we show the metric mean dimension of the product of the self-similar measure is closely related to the similarity dimension of IFS.
Let $m\geq 1$ be a positive integer and endow $\mathbb{R}^{m}$ with the supremum norm $\|\cdot\|_{\mathbb{R}^{m}}$. Assume that $T_{i}:\mathbb{R}^{m}\to\mathbb{R}^{m}$ is contracting similitude for $1\leq i\leq l$, i.e., for some $r_{i}\in(0,1)$,
$\|T_{i}x-T_{i}y\|_{\mathbb{R}^{m}}=r_{i}\|x-y\|_{\mathbb{R}^{m}}.$
The iterated function system (IFS for short) $\{T_{i}\}_{1\leq i\leq l}$ satisfies the Open Set Condition if there exists a bounded open subset $V\subset\mathbb{R}^{m}$ such that $T_{i}(V)\subset V,\ \forall 1\leq i\leq l$, and $T_{i}(V)\cap T_{j}(V)=\emptyset~\text{for}~i\neq j.$ For instance, the middle-third Cantor set, the Sierpinński triangle or gasket, are the attractors of iterated function systems satisfying the Open Set Condition. Under the OSC condition, it is well known that there are (unique) compact subset $K$, which we call the self-similar set, and a measure $\mu$, which we call the self-similar measure, such that
@@
    ${\rm dim}_{H}(K,d)={\rm dim}_{B}(K,d)=\alpha$, and $\alpha$ is the unique solution of equation
@@
    where ${\rm dim}_{H}$ denotes the Hausdorff dimension under the metric $d:=\|\cdot\|_{\mathbb{R}^{m}}$. The root $\alpha$ is called the similarity dimension of the IFS;
@@
    $\mu=\sum_{i=1}^{l}r_{i}^{\alpha}\cdot\mu\circ T_{i}^{-1}$ and ${\rm supp}(\mu)=K$.
Now take the attractor $K$ obtained as above and the product measure $\nu:=\mu^{\otimes\mathbb{Z}}$ on $K^{\mathbb{Z}}$.
@@
The three types of metric mean dimensions of $G_{\nu}$ coincide with the similarity dimension. i.e.,
$\overline{\rm mdim}_{M}^{S}(\sigma,G_{\nu},d^{\mathbb{Z}})=\alpha=\limsup_{\epsilon\to 0}\frac{\inf_{{\rm diam}(\xi)<\epsilon}h_{\nu}(\sigma,\xi)}{|\log\epsilon|},$
where $S\in\{B,P,UC\}$.
@@
In [Hut81], Hutchinson proved that there exists a constant $C>0$ such that for any $z\in K$ and $\epsilon\in(0,1)$, one has
$\mu(B(z,\epsilon))\leq C\epsilon^{\alpha}.$ (4.5)
Fix $\epsilon\in(0,1)$. For any $x=(x_{k})_{k\in\mathbb{Z}}\in K^{\mathbb{Z}}$, we define
$L_{n}(x,\epsilon)=\left\{(y_{k})_{k\in\mathbb{Z}}\in K^{\mathbb{Z}}:\|x_{k}-y_{k}\|_{\mathbb{R}^{m}}<\epsilon,\ 0\leq k\leq n-1\right\}.$
Then $B_{n}(x,\epsilon)\subset L_{n}(x,\epsilon)$. Since $\nu(L_{n}(x,\epsilon))\leq C^{n}\epsilon^{n\alpha}$ by (4.5), one has
@@
which yields that the Brin-Katok local $\epsilon$-entropy of $\nu$ is bounded blew by
$\overline{h}_{\nu}^{BK}(\sigma,\epsilon):=\int\limsup_{n\to\infty}\frac{-\log\mu(B_{n}(x,\epsilon))}{n}d\nu\geq\alpha|\log\epsilon|-\log C.$
Applying the Shannon-McMillan-Breiman theorem, the following inequality is valid for any TDS $(X,d,f)$: